# Nonparametric Uniform Inference in

### v1

@@ 5.2 Large Sample Theory for Surrogate Policy Function Estimation
    (Full-Sample) For each $\ell\in[m]$ and any $g_{1},g_{2}\in\mathcal{S}_{n}$ of $g_{\ast}$, the full-sample nuisance parameter estimator satisfies
@@
Finally, we impose a smoothness condition on the surrogate loss and the surrogate policy function to facilitate the derivation of their asymptotic properties.
@@
Assumption 5.4 (Surrogate Policy):
The surrogate policy function $g_{\ast}$ is uniformly bounded on compact $\mathcal{X}$ and belongs to a Hölder ball $\Lambda^{h}(L)$ of smoothness order $h>1/2$ and radius $L\in(0,\infty)$,  [^9] See (B.1) in the appendix for the definition of the Hölder ball $\Lambda^{h}(L)$ of smoothness order $h$ and radius $L$.
For standardization, we introduce the variance process $\sigma(x)^{2}=p_{1:k}(x)^{\prime}Q^{-1}\Sigma Q^{-1}p_{1:k}(x)$, where
@@
We may estimate it by the sample counterpart $\widehat{\sigma}(x)^{2}=p_{1:k}(x)^{\prime}\widehat{Q}^{-1}\widehat{\Sigma}\widehat{Q}^{-1}p_{1:k}(x)$ where
@@
With the above definitions and notation in place, the following theorem establishes the uniform convergence rate and Gaussian coupling for the surrogate policy function estimator $\widehat{g}$.
@@
5.3 Bootstrap Inference for the Surrogate Policy Function
@@
To compute the critical value $cv_{n}^{b}(1-\alpha)$ of the supremum statistic $\sup_{x\in\mathcal{X}}|t(x)|$ in (5.13), one can evaluate $\sup_{x\in\mathcal{X}}|t^{b}_{g}(x)|$ based on a large number of independent draws of $\{\omega_{i}\}_{i}$. The critical value $cv_{n}^{b}(1-\alpha)$ is then approximated as
@@
Construct the uniform confidence band (UCB) for the surrogate policy function $g_{\ast}$ as
@@
To compute the one-sided critical values, one can evaluate $\sup_{x\in\mathcal{X}}t^{b}_{g}(x)$ and $\inf_{x\in\mathcal{X}}t^{b}_{g}(x)$ based on a large number of independent draws of $\{\omega_{i}\}_{i}$, as before. The upper and lower critical values, $cv_{n}^{1,b}(1-\alpha)$ and $cv_{n}^{1,b}(\alpha)$, are then approximated as
@@ 6 Estimation and Inference for the Optimal Policy Value
$V(g)$ $:=\mathbb{E}\!\left[\psi_{1}(Z_{i},\eta)\,\mathbf{1}\{g(X_{i})\geq 0\}+\psi_{0}(Z_{i},\eta)\,\mathbf{1}\{g(X_{i})<0\}\right],$
@@
$V(g)=\mathbb{E}[\overline{\psi}(Z_{i},\eta)]-L(g),$
where $L$ is defined in (2.1). Therefore, the surrogate policy $g_{\ast}$ identified via the surrogate loss approach attains the maximum level of welfare within $\mathcal{G}$; that is, $V_{0}=V(g_{\ast})$. Although the surrogate policy $g_{\ast}$ is obtained by minimizing the surrogate risk $L^{\ast}$ in (4.1) to ensure point identification, this surrogate risk $L^{\ast}$ does not coincide with the value of the original risk $L$ in (2.1). To estimate the optimal value $V_{0}$, we need to evaluate the original risk $L$, rather than the surrogate risk $L^{\ast}$, at an estimate of $g_{\ast}$.
@@ 6.1 Estimation of the the Optimal Policy Value
Let $\widehat{g}$ denote a preliminary estimator of the surrogate decision rule $g_{\ast}$, a specific example of which is provided in Section 5.1. In general, $\widehat{g}$ may be any estimator that satisfies a mild convergence rate condition, to be formally stated as Assumption 6.1 shortly. Accordingly, Sections 6.16.3 proceed with an arbitrary estimator $\widehat{g}$, whereas Section 6.4 verifies that our particular estimator from Section 5 is a viable candidate.
To account for the estimation of $\eta$, we employ a cross-fitting procedure to estimate the optimal policy value. Partition $n]=\{1,\ldots,n\}$ into $m$ folds $I_{1},\ldots,I_{m}$. Assume that the folds are approximately balanced, i.e., $n_{(1)}\asymp\cdots\asymp n_{(m)}$, where $n_{(\ell)}:=|I_{(\ell)}|$, and that $m$ is fixed. For a given fold $\ell\in[m]=\{1,\ldots,m\}$, obtain an estimator $\widehat{\eta}_{(\ell)}$ of $\eta_{0}$ using the complementary subsample $I_{(\ell)}^{c}$. We then estimate the optimal policy value $V_{0}$ via the plug-in cross-fitted estimator
$\widehat{V}_{n}(\widehat{g}):=\frac{1}{n}\sum_{\ell=1}^{m}\sum_{i\in I_{(\ell)}}\Big[\widehat{\psi}_{1,(\ell)}(Z_{i})\,\mathbf{1}\{-\widehat{g}(X_{i})<0\}+\widehat{\psi}_{0,(\ell)}(Z_{i})\,\mathbf{1}\{\widehat{g}(X_{i})<0\}\Big],$ (6.1)
@@ 6.2 Asymptotic Properties for the Optimal Policy Value
The optimal policy value estimator in (6.1) is constructed via a plug-in rule with a preliminary estimator $\widehat{g}$ of the surrogate policy $g_{\ast}$. For valid inference, $\widehat{g}$ must satisfy conditions ensuring that both the estimation error and the approximation error are dominated and hence do not affect the limiting distribution. At first glance, this requirement may appear stringent; however, owing to the optimality of $g_{\ast}$ for $V$, it is in fact attainable under the following mild condition.
Assumption 6.1 (Policy Function Estimator):
Let $g_{k}=\arg\min_{g\in\mathcal{G}_{k}}\sup_{x\in\mathcal{X}}|g_{\ast}(x)-g(x)|$ and $B_{k}:=\sup_{x\in\mathcal{X}}|g_{\ast}(x)-g_{k}(x)|$. A preliminary estimator $\widehat{g}$ and the associated projection $g_{k}$ satisfy
$\|\widehat{g}-g_{k}\|_{\mathbb{P},\infty}=\sup_{x\in\mathcal{X}}|\widehat{g}(x)-g_{k}(x)|=o_{p}(n^{-1/4})\qquad\text{and}\qquad B_{k}=o(n^{-1/2}).$
Assumption 6.1 is stated in terms of the optimizer $g_{\ast}$ of the surrogate loss for internal consistency of our theoretical framework. As will be formally shown in Section 6.4, this high-level condition is satisfied in particular by our cross-fitted sieve estimator proposed in Section 5. However, our procedure can also accommodate a broad class of estimators that converge to some $g_{\ast\ast}$, satisfying $V(g_{\ast\ast})=V_{0}$, at a rate faster than $o_{p}(n^{-1/4})$. Examples include $\ell_{1}$-penalized estimators (Athey and Wager, 2021), parametric methods based on the conventional 0-11 loss (e.g., Kitagawa and Tetenov, 2018), and the nonparametric CATE estimator (e.g., Armstrong and Shen, 2023), as long as these estimators converges to $\mathbb{E}[\psi_{1}(Z,\eta)|X=x]-\mathbb{E}[\psi_{0}(Z,\eta)|X=x]$ uniformly over $\mathcal{X}$ at a rate faster than $o_{p}(n^{-1/4})$. Our approach therefore provides a flexible limit theory and ensuring asymptotically valid inference for a range of policy estimators from the literature.
While we can accommodate a general class of weight functions $\psi_{1}$ and $\psi_{0}$, the following condition is imposed to eliminate any possible irregular behaviors.
@@
The random variable $U_{i}:=\mathbb{E}[\psi_{1}(Z_{i},\eta)|X_{i}]-\mathbb{E}[\psi_{0}(Z_{i},\eta)|X_{i}]$ admits a continuous probability density function $f_{U}$ over a small neighborhood $\delta,\delta]$ of zero.
Assumption 6.2 corresponds to the standard margin condition widely used in the literature (see, for example, Kitagawa and Tetenov (2018) and Ponomarev and Semenova (2024)). This assumption requires the data to exhibit a “low-noise” property. Specifically, if the random variable $U_{i}$ has a continuous density around zero, then $\mathbb{P}(|U_{i}|<t)=O(t)$ for sufficiently small $t>0$, which further implies $\mathbb{E}[|U_{i}|1\{|U_{i}|<t\}]=O(t^{2})$. Intuitively, when $|U_{i}|$ is small—meaning that the welfare difference between the two actions is nearly tied—the estimated optimal rule is more likely to misclassify the optimal decision, but each such misclassification incurs only a small welfare loss. If these near-tie events occur infrequently, as ensured by the margin condition, the expected welfare loss regarding the optimal rule is even smaller. So, the margin condition ensures that the population welfare functional $V(g)$ is not too sensitive to small deviations of $g$ from the optimal policy $g_{\ast}$. [^10] Assumption 6.2 corresponds to the margin condition in Kitagawa and Tetenov (2018), where $\mathbb{P}(|U_{i}|<t)=O(t^{\alpha})$ with $\alpha=1$. Our results remain valid under any distribution of $U_{i}$ satisfying $\mathbb{P}(|U_{i}|<t)=O(t^{\alpha})$ for $\alpha\geq 1$, but generally fail when $\alpha<1$.
@@
    (Boundedness) For $j=0,1$, $\mathbb{E}[|\psi_{j}(Z_{i},\eta)|^{q}]$ is bounded uniformly in $\eta$ for some $q>4$, and $\mathbb{E}[\psi_{j}(Z_{i},\eta)|X_{i}=x]$ is bounded over the support of $X$.
Assumption 6.3.(i) requires that the cross-fitted estimator $\widehat{\eta}$ of the nuisance parameter induces asymptotically negligible estimation errors. This condition has been also employed by Athey and Wager (2021, Assumption 2) in the context of Example III, and can be achieved by employing a Neyman-orthogonal score, which ensures that first-step estimation errors are asymptotically negligible provided that $\widehat{\eta}_{(\ell)}$ converges at the $n^{-1/4}$ rate, rather than the parametric $n^{-1/2}$ rate. Such convergence rates are attainable with a variety of machine learning methods. For example, the Lasso achieves the $n^{-1/4}$ rate under approximate sparsity and restricted eigenvalue conditions (Bickel et al., 2009; Belloni et al., 2015). Random forests and boosted trees are valid provided the regression function is sufficiently smooth and sample splitting is employed (Bühlmann and Yu, 2002; Wager and Athey, 2018). Deep neural networks with controlled depth and bounded weight norms can approximate smooth functions at nearly minimax rates (Chen and White, 1999; Yarotsky, 2017; Schmidt-Hieber, 2020), among others. Assumption 6.3.(ii) further imposes uniform bounds on the weight functions, $\psi_{1}$ and $\psi_{0}$. Note that this condition allows for unbounded outcome variable $Y$ in the welfare maximization (Example III) and is also satisfied in the expected utility maximization with binary action (Example II) under the Condition 1 of Elliott and Lieli (2013). It holds trivially in the maximum score estimation (Example I).
Under these rate conditions and the boundedness conditions, we obtain the following asymptotic normality result for the plug-in estimator of the optimal policy value.
@@
In addition to the conditions for Theorem 4.1(ii), suppose that Assumptions 5.3 and Assumptions 6.1 to 6.3 hold. For any sequence of data generating processes $\mathbb{P}_{n}$, we have
@@
While the proof in the appendix formally establishes the results of Theorem 6.1, we provide the intuition here, particularly illustrating why the estimation effect of the preliminary estimator $\widehat{g}$ that satisfies Assumption 6.1 is asymptotically negligible despite the fact that it appears in the indicator function and converges at a slower rate than the root-$n$ rate. Consider the decomposition:
@@
Hirano and Porter (2012, Section 2.4) highlight that nondifferentiability in a population criterion function prevents regular estimation. In particular, such an impossibility phenomenon arises in estimating $\phi(\theta)=\max\{\theta_{1},\theta_{2}\}$, which is non-differentiable at the kink point $\theta_{1}=\theta_{2}$. This kink causes the plug-in estimator $\max\{\widehat{\theta}_{1},\widehat{\theta}_{2}\}$ to have a non-Gaussian distribution when $\theta_{1}=\theta_{2}$, even if $(\widehat{\theta}_{1},\widehat{\theta}_{2})$ were Gaussian and centered at $(\theta_{1},\theta_{2})$.
@@
$\sup_{g\in\mathcal{G}}V(g)$ $=\sup_{g\in\mathcal{G}}\mathbb{E}\!\left[\mathbb{E}[\psi_{1}(Z_{i},\eta)\mid X_{i}]\cdot\mathbf{1}\{g(X_{i})\geq 0\}+\mathbb{E}[\psi_{0}(Z_{i},\eta)\mid X_{i}]\cdot\mathbf{1}\{g(X_{i})<0\}\right]$
@@
rather than $\max\{\theta_{1},\theta_{2}\}$, where $\vartheta_{1}:=\mathbb{E}[\psi_{1}(Z_{i},\eta)\mid X_{i}]$ and $\vartheta_{2}:=\mathbb{E}[\psi_{0}(Z_{i},\eta)\mid X_{i}]$ are random variables. Hence, the two problems differ mathematically. As emphasized in Remark 1, given the smoothness of the population objective, asymptotic normality is attainable even with a non-smooth sample criterion, with the first-order condition for the optimality of $g_{\ast}$ playing the crucial role. $\blacktriangle$

### v2

@@ 5.2 Large Sample Theory for the Representative Policy Function Estimation
    (Full-Sample) For any $g_{1},g_{2}\in\mathcal{S}_{n}$ of $g_{\ast}$, the full-sample nuisance parameter estimator satisfies
@@
Finally, we impose a smoothness condition on the surrogate loss and the representative policy function to facilitate the derivation of their asymptotic properties.
@@
Assumption 5.4 (Representative Policy):
The representative classification policy function $g_{\ast}$ is uniformly bounded on compact $\mathcal{X}$ and belongs to a Hölder ball $\Lambda^{h}(L)$ of smoothness order $h>1/2$ and radius $L\in(0,\infty)$. [^9] See (B.1) in the appendix for the definition of the Hölder ball $\Lambda^{h}(L)$ of smoothness order $h$ and radius $L$.
For standardization, we introduce the variance process $\sigma^{2}(x)=p_{1:k}(x)^{\prime}Q^{-1}\Sigma Q^{-1}p_{1:k}(x)$, where
@@
We may estimate it by the sample counterpart $\widehat{\sigma}^{2}(x)=p_{1:k}(x)^{\prime}\widehat{Q}^{-1}\widehat{\Sigma}\widehat{Q}^{-1}p_{1:k}(x)$ where
@@
With the above definitions and notation in place, the following theorem establishes the uniform convergence rate and Gaussian coupling for the representative policy function estimator $\widehat{g}$.
@@
5.3 Bootstrap Inference for the Representative Policy Function
@@
To compute the critical value $cv_{n}(1-\alpha)$ of the supremum statistic $\sup_{x\in\mathcal{X}}|t(x)|$ in (5.13), one can evaluate $\sup_{x\in\mathcal{X}}|t^{b}_{g}(x)|$ based on a large number of independent draws of $\{\omega_{i}\}_{i}$. The critical value $cv_{n}(1-\alpha)$ is then approximated as
@@
Construct the uniform confidence band (UCB) for the representative policy function $g_{\ast}$ as
@@
To compute the one-sided critical values, one can evaluate $\sup_{x\in\mathcal{X}}t^{b}_{g}(x)$ and $\inf_{x\in\mathcal{X}}t^{b}_{g}(x)$ based on a large number of independent draws of $\{\omega_{i}\}_{i}$, as before. The upper and lower critical values, $cv_{n}^{1,b}(1-\alpha)$ and $cv_{n}^{1,b}(\alpha)$, are then computed as
@@ 6 Estimation and Inference for the Optimal Policy Value
$V(g)$ $:=\mathbb{E}\!\left[\psi_{1}(Z_{i},\eta_{0})\,\mathbf{1}\{g(X_{i})\geq 0\}+\psi_{0}(Z_{i},\eta_{0})\,\mathbf{1}\{g(X_{i})<0\}\right],$
@@
$V(g)=\mathbb{E}[\overline{\psi}(Z_{i},\eta_{0})]-L(g),$
where $L$ is defined in (2.1). Therefore, the representative classification policy function $g_{\ast}$ identified via the surrogate loss approach attains the maximum level of welfare within $\mathcal{G}$; that is, $V_{0}=V(g_{\ast})$. Although the representative policy function $g_{\ast}$ is obtained by minimizing the surrogate risk $L^{\ast}$ in (4.1) to ensure point identification, this surrogate risk $L^{\ast}$ does not coincide with the value of the original risk $L$ in (2.1). To estimate the optimal value $V_{0}$, we need to evaluate the original risk $L$, rather than the surrogate risk $L^{\ast}$, at an estimate of $g_{\ast}$.
@@ 6.1 Estimation of the the Optimal Policy Value
Let $\widehat{g}$ denote the estimator of the representative classification policy function $g_{\ast}$, as obtained in Section 5.1. To account for the estimation of $\eta$, we employ a cross-fitting procedure to estimate the optimal policy value. Partition $n]=\{1,\ldots,n\}$ into $m$ folds $I_{1},\ldots,I_{m}$. Assume that the folds are approximately balanced, i.e., $n_{(1)}\asymp\cdots\asymp n_{(m)}$, where $n_{(\ell)}:=|I_{(\ell)}|$, and that $m$ is fixed. For a given fold $\ell\in[m]=\{1,\ldots,m\}$, obtain an estimator $\widehat{\eta}_{(\ell)}$ of $\eta_{0}$ using the complementary subsample $I_{(\ell)}^{c}$. We then estimate the optimal policy value $V_{0}$ via the plug-in cross-fitted estimator
$\widehat{V}_{n}(\widehat{g}):=\frac{1}{n}\sum_{\ell=1}^{m}\sum_{i\in I_{(\ell)}}\Big[\widehat{\psi}_{1,(\ell)}(Z_{i})\,\mathbf{1}\{-\widehat{g}(X_{i})\leq 0\}+\widehat{\psi}_{0,(\ell)}(Z_{i})\,\mathbf{1}\{\widehat{g}(X_{i})<0\}\Big],$ (6.1)
@@ 6.2 Asymptotic Properties for the Optimal Policy Value
The optimal policy value estimator (6.1) is constructed via a plug-in rule using our estimator $\widehat{g}$ of the representative policy function $g_{\ast}$, as given in Section 5.1. For valid inference, $\widehat{g}$ must satisfy conditions ensuring that both its estimation error and its surrogate misspecification error are dominated and hence do not affect the limiting distribution of the policy value estimator. Theorem 5.1 shows that our estimator $\widehat{g}$ converges at the rate $\xi_{k}\sqrt{\log(k)/n}$. We verify below that, under a set of mild conditions, this nonparametric rate suffices to ensure its impact on the distribution of the policy value estimator is asymptotically negligible.
@@
The random variable $U_{i}:=\mathbb{E}[\psi_{1}(Z_{i},\eta_{0})|X_{i}]-\mathbb{E}[\psi_{0}(Z_{i},\eta_{0})|X_{i}]$ admits a continuous probability density function $f_{U}$ over a small neighborhood $\delta,\delta]$ of zero.
Assumption 6.1 corresponds to the standard margin condition widely used in the empirical welfare maximization literature (see, for example, Kitagawa and Tetenov (2018) and Ponomarev and Semenova (2024)). This assumption requires the data to exhibit a “low-noise” property. Specifically, if the random variable $U_{i}$ has a continuous density around zero, then $\mathbb{P}(|U_{i}|<t)=O(t)$ for sufficiently small $t>0$, which further implies $\mathbb{E}[|U_{i}|1\{|U_{i}|<t\}]=O(t^{2})$. Intuitively, when $|U_{i}|$ is small—meaning that the welfare difference between the two actions is nearly tied—the estimated optimal rule is more likely to misclassify the optimal decision, but each such misclassification incurs only a small welfare loss. If these near-tie events occur infrequently, as ensured by the margin condition, the expected welfare loss regarding the optimal rule is even smaller. So, the margin condition ensures that the population welfare functional $V(g)$ is not too sensitive to small deviations of $g$ from the optimal policy $g_{\ast}$. [^10] Assumption 6.1 corresponds to the margin condition in Kitagawa and Tetenov (2018), where $\mathbb{P}(|U_{i}|<t)=O(t^{\alpha})$ with $\alpha=1$. Our results remain valid under any distribution of $U_{i}$ satisfying $\mathbb{P}(|U_{i}|<t)=O(t^{\alpha})$ for $\alpha\geq 1$, but generally fail when $\alpha<1$.
@@
    (Boundedness) For $j=0,1$, and uniformly over the shrinking neighborhood $\mathcal{T}_{n}$ of $\eta_{0}$$\mathbb{E}[|\psi_{j}(Z_{i},\eta)|^{q}]$ is bounded for some $q>4$, and $\mathbb{E}[\psi_{j}(Z_{i},\eta)|X_{i}=x]$ is bounded over the support of $X$.
Assumption 6.2.(i) requires that the cross-fitted estimator $\widehat{\eta}$ of the nuisance parameter induces asymptotically negligible estimation errors. This condition has been also employed by Athey and Wager (2021, Assumption 2) in the context of Example III, and can be achieved by employing a Neyman-orthogonal score, which ensures that first-step estimation errors are asymptotically negligible provided that $\widehat{\eta}_{(\ell)}$ converges at the $n^{-1/4}$ rate, rather than the parametric $n^{-1/2}$ rate. Such convergence rates are attainable with a variety of machine learning methods. For example, the Lasso achieves the $n^{-1/4}$ rate under approximate sparsity and restricted eigenvalue conditions (Bickel et al., 2009; Belloni et al., 2015). Random forests and boosted trees are valid provided the regression function is sufficiently smooth and sample splitting is employed (Bühlmann and Yu, 2002; Wager and Athey, 2018). Deep neural networks with controlled depth and bounded weight norms can approximate smooth functions at nearly minimax rates (Chen and White, 1999; Yarotsky, 2017; Schmidt-Hieber, 2020), among others. Assumption 6.2.(ii) further imposes uniform bounds on the (conditional) expectations of weight functions, $\psi_{1}$ and $\psi_{0}$. Note that this condition allows for unbounded outcome variable $Y$ in the welfare maximization (Example III) and is also satisfied in the expected utility maximization with binary action (Example II) under the Condition 1 of Elliott and Lieli (2013). It holds trivially in the maximum score estimation (Example I).
Under these rate conditions and the boundedness conditions, we obtain the following asymptotic normality result for our estimator of the optimal policy value.
@@
In addition to the conditions for Theorem 5.1, suppose that Assumptions 6.1 and 6.2 hold, and the sieve order $k$ satisfies $k^{4+\varepsilon}/n\lesssim 1$ for some $\varepsilon>0$. Then, for any sequence of data generating processes $\mathbb{P}_{n}$, we have
@@
While the proof in the appendix formally establishes the results of Theorem 6.1, we provide the intuition here, particularly illustrating why the estimation effect of our classification policy function estimator $\widehat{g}$ as given in Section 5.1 is asymptotically negligible despite the fact that it appears in the indicator function and converges at a slower rate than the root-$n$ rate. Consider the decomposition:
@@
Although the plug-in value estimator in (6.1) is constructed using the estimator $\widehat{g}$ described in Section 5.1, our inferential results do not rely on the specific form of this estimator. What is required for valid second-stage inference is that the first-stage policy function estimator converges sufficiently fast so that its estimation error becomes asymptotically negligible. Indeed, with slight modifications on the proof, we can show that any estimator $\widetilde{g}$ would be admissible if it converges uniformly over $\mathcal{X}$ to a limit $g_{\ast\ast}$ satisfying $V(g_{\ast\ast})=V_{0}$ at a rate faster than $o_{p}(n^{-1/4})$. That is, our framework accommodates a broad class of first-stage methods, including $\ell_{1}$-penalized estimators (Athey and Wager, 2021), parametric approaches based on the 0–1 loss (e.g., Kitagawa and Tetenov, 2018), and nonparametric CATE-based methods (e.g., Bhattacharya and Dupas, 2012; Armstrong and Shen, 2023; Park, 2025), provided they achieve the uniform convergence rate $o_{p}(n^{-1/4})$ to $\mathbb{E}[\psi_{1}(Z,\eta_{0})\mid X=x]-\mathbb{E}[\psi_{0}(Z,\eta_{0})\mid X=x]$. Thus, while our construction employs the estimator $\widehat{g}$ for the representative classification policy $g_{\ast}$, the resulting limit theory supports valid inference for a wide range of first-stage estimators used in the policy learning literature.
Remark 2:
@@
Hirano and Porter (2012, Section 2.4) highlight that non-differentiability in a population criterion function prevents regular estimation. In particular, such an impossibility phenomenon arises in estimating $\phi(\theta)=\max\{\theta_{1},\theta_{2}\}$, which is non-differentiable at the kink point $\theta_{1}=\theta_{2}$. This kink causes the plug-in estimator $\max\{\widehat{\theta}_{1},\widehat{\theta}_{2}\}$ to have a non-Gaussian distribution when $\theta_{1}=\theta_{2}$, even if $(\widehat{\theta}_{1},\widehat{\theta}_{2})$ were Gaussian and centered at $(\theta_{1},\theta_{2})$.
@@
$\sup_{g\in\mathcal{G}}V(g)$ $=\sup_{g\in\mathcal{G}}\mathbb{E}\!\left[\mathbb{E}[\psi_{1}(Z_{i},\eta_{0})\mid X_{i}]\cdot\mathbf{1}\{g(X_{i})\geq 0\}+\mathbb{E}[\psi_{0}(Z_{i},\eta_{0})\mid X_{i}]\cdot\mathbf{1}\{g(X_{i})<0\}\right]$
@@
rather than $\max\{\theta_{1},\theta_{2}\}$, where $\vartheta_{1}:=\mathbb{E}[\psi_{1}(Z_{i},\eta_{0})\mid X_{i}]$ and $\vartheta_{2}:=\mathbb{E}[\psi_{0}(Z_{i},\eta_{0})\mid X_{i}]$ are random variables. Hence, the two problems differ mathematically. As emphasized in Remark 2, given the smoothness of the population objective, asymptotic normality is attainable even with a non-smooth sample criterion, with the first-order condition for the optimality of $g_{\ast}$ playing the crucial role. $\blacktriangle$